OpenStudy (anonymous):

A square with the side w is drawn inside a 6-unit square. What equation represents the area of the 6-unit square excluding the area of the square inside it.

hartnn (hartnn):

 area of 6 unit square = 6*6 = 36 area of w unit square =w*w areas of required shaded region = 36-w*w